OpenStudy (ikou):

Please help me with this! Two mechanics worked on a car. The first mechanic worked for 5 hours, and the second mechanic worked for 15 hours. Together they charged a total of $1950. What was the rate charged per hour by each mechanic if the sum of the two rates was $200 per hour?

OpenStudy (mathstudent55):

Let's call the rates x and y. If the sum of the rates is $200, can you write an equation using x and y to represent that fact?

OpenStudy (ikou):

Well I know that there can be two equations the Value eq: 5x+15y+1950 and the Rate eq: x+y=200

OpenStudy (mathstudent55):

Yes. I was asking for the rate equation only, but you have both equations already, and they are correct. Now you need to solve the system of equations.

OpenStudy (ikou):

I don't really know how though

OpenStudy (mathstudent55):

Ok. I'll explain it to you. One method of solving a system of equation is called "the substitution method." Let's use that method here.

OpenStudy (mathstudent55):

The first step is to solve one equation for one variable. We see that is easier to do that with the rate equation. Can you solve the rate equation, x + y = 200, for x?

OpenStudy (ikou):

what happens to the y, and do I plug in 5x where the x should be

OpenStudy (ikou):

so if I plug in 5x to the equation and just ignore the y then I'd get 50, so x earns 50 an hour

OpenStudy (mathstudent55):

First just solve x + y = 200 for x. What do you get?

OpenStudy (mathstudent55):